OpenStudy (samigupta8):

a particle of mass m is tied to a light string of length l and roated along a vertical circular pathh.What should be the minimum speed at the highest point of its path so that the string does not become slack at any position?

OpenStudy (anonymous):

The centrifugal force, act upon mass, must be grater than weight at the top of the flight.

OpenStudy (samigupta8):

no but its options are as follows....\[a.\sqrt{2gl} \]\[b.\sqrt{gl}\]\[c.0\]\[d.\sqrt{gl/2} \]

OpenStudy (anonymous):

so, \[\frac{ mv ^{2} }{ l } \ge mg\] if the centrifugal force equals weight, then there is not any tension on the string.. you can take them as equal, but be careful what i mention above.

OpenStudy (samigupta8):

plss...cn u explain it with help of a suitable diagram

OpenStudy (anonymous):

 Centrifugal force - weight = tension